OpenStudy (anonymous):

Solve the problem by using substitution. Be sure to show all your work. x+y=5 y=x-1

OpenStudy (across):

1. Substitute \(y=x-1\) into \(y+x=5\):\[x+(x-1)=5,\\2x-1=5,\\x=3.\]Substitute \(x=3\) into \(y=x-1\):\[y=3-1=2.\]